OpenStudy (anonymous):

A can do a piece of work in 24 days. When he had worked for 4 days B joined him, if the work was completed in 16 days, in how many days B alone can finish the work?

OpenStudy (anonymous):

How much A can do in 1 day?

OpenStudy (anonymous):

work/24

OpenStudy (anonymous):

Good, Let B and do the work in x days, then solve this equation for x \[ \frac{1}{6}+12 \left(\frac{1}{24}+\frac{1}{x}\right)=1\] x= 36 days

OpenStudy (anonymous):

I GET 90

OpenStudy (anonymous):

Do you know the right answer?

OpenStudy (anonymous):

I DON4T KnOW IF they compleet it after or in 16days !!!! maybe your true th

OpenStudy (anonymous):

acctualy the right unsewer is 96

OpenStudy (anonymous):

ow so the 4 days is not included in those 16 days. So in my equation you have to multiply by 16 instead of 12.